OpenStudy (anonymous):

can anyone help me multiply and simplify this: sqrt7x(sqrtx -7 sqrt7)

OpenStudy (anonymous):

\[\sqrt{7x}(\sqrt{x}-7\sqrt{7})\]

OpenStudy (anonymous):

yes. can you show me how to do it?